### v1

# mystyle backgroundcolor=white!10

Abstract
This paper studies finite-time safety and reach-avoid verification for stochastic discrete-time dynamical systems. The aim is to ascertain lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). We introduce novel barrier-like sufficient conditions for characterizing these bounds, which either complement existing ones or fill gaps. Finally, we demonstrate the efficacy of these conditions on two examples.
Index Terms:
Finite-time Safety/Reach-avoid Verification; Probability Bounds; Stochastic Discrete-time Systems
I Introduction
Temporal verification plays a pivotal role in modern systems analysis, especially in the realm of complex systems where temporal behavior holds utmost significance [10]. It entails a rigorous scrutiny of a system’s compliance with temporal properties, including safety and reach-avoid guarantees, to ensure desired outcomes while circumventing undesirable events. Formal methods such as model checking [4, 1] and theorem proving [7] serve as indispensable tools in this endeavor, enabling precise and thorough analysis of temporal specifications.
Barrier certificates were initially proposed for deterministic systems as a popular formal approach to temporal verification in [8]. They offer Lyapunov-like assurances regarding system behavior, with the mere existence of a barrier function being sufficient to establish the satisfiability of safety and/or reachability specifications in [10]. Subsequent efforts have focused on adapting and enhancing deterministic barrier functions, as well as broadening their applications [2, 14]. However, many real-world applications are susceptible to stochastic disturbances and are thus modeled as stochastic systems. In the stochastic setting, safety verification over the infinite time horizon via barrier certificates was introduced alongside its deterministic counterpart in [9]. Utilizing Ville’s Inequality [12], [9] constructed a non-negative barrier function and provided a sufficient condition for upper bounding the probability of eventually entering an unsafe region from a given initial state while remaining within a state constraint set. More recently, a new barrier function, constructed by relaxing a set of equations, was proposed for lower bounding the probability of eventually entering an unsafe or desired region from an initial state while adhering to state constraints in [13, 15]. These barrier functions were further extended to determine the lower and upper bounds of the liveness probability over the infinite time horizon for a specified safe set and set of initial states in [16]. Notably, these proposed barrier functions require the expected value evolution to be monotonically non-increasing or non-decreasing, which can be demanding and thus limit their practical applications. On the other hand, finite-time temporal verification holds greater practical significance as most real-world systems operate within well-defined time constraints. Drawing inspiration from [6], the concept of a $c$-martingale was introduced in [11] for stochastic continuous-time systems modelled by stochastic differential equations, enabling a controlled increase in the expected certificate value at each time step and offering an upper bound on the probability of reaching unsafe sets within bounded time horizons. Building upon this, [5] extended the $c$-martingale framework to address temporal logic verification for discrete-time systems, further extending the applicability of this approach.
This paper investigates the finite-time safety and reach-avoid verification of stochastic discrete-time systems. The finite-time safety verification problem aims to compute both lower and upper bounds of the probability that a system, starting from an initial state in a safe set, will exit the safe set throughout a given bounded time interval. From a reachability perspective, it involves computing lower and upper bounds on the probability of reaching the complement of the safe set within the specified bounded time interval. Thus, it exclusively addresses safety or reachability concerns. In contrast, finite-time reach-avoid verification integrates guarantees of safety and reachability. It seeks to establish lower and upper bounds on the probability that a system, starting from an initial state in a safe set, will reach a target set within a designated bounded time interval while ensuring it remains within the safe set before reaching the target set. Although these two problems are interconnected, they differ fundamentally in essence, as we will elaborate on in the preliminaries section. We propose novel barrier-like conditions to address these two problems. These conditions either complement existing ones or fill gaps, facilitating the attainment of tight probability bounds. Finally, we demonstrate the effectiveness of the proposed conditions on two numerical examples, utilizing semi-definite programming tools.
The main contributions of this work are summarized below.
  1. 1.
    This work studies the finite-time safety and reach-avoid verification of stochastic discrete-time systems. Compared with existing works [5, 6], which merely offered upper bounds on the relevant probabilities, our contribution goes beyond by providing both lower and upper bounds. These bounds deepen our understanding of the system’s characteristics and yield a more accurate estimation of the true probability, thus enhancing the overall rigour and precision of the finite-time safety and reach-avoid analysis.
  2. 2.
    The proposed barrier-like conditions for upper bounding the probabilities complement existing ones in [5, 6], facilitating the gain of sharper upper bounds.
This paper is structured as follows: in Section II, we formulate the finite-time safety and reach-avoid verification problems of interest. In Section III and IV, we introduce our barrier-like conditions for addressing these two problems, respectively. After demonstrating the performance of the proposed conditions on two examples in Section V, we conclude this paper in Section VI.
II Preliminaries
We start the exposition by a formal introduction of discrete-time systems subject to stochastic disturbances and finite-time safety/reach-avoid verification problems of interest. Before posing the problem studied, let me introduce some basic notions used throughout this paper: $\mathbb{R}$ denotes the set of real values; $\mathbb{N}$ denotes the set of nonnegative integers; $\mathbb{N}_{\leq k}$ is the set of non-negative integers being less than or equal to $k$; for sets $\Delta_{1}$ and $\Delta_{2}$, $\Delta_{1}\setminus\Delta_{2}$ denotes the difference of sets $\Delta_{1}$ and $\Delta_{2}$, which is the set of all elements in $\Delta_{1}$ that are not in $\Delta_{2}$; $1_{A}(\bm{x})$ denotes the indicator function in the set $A$, where, if $\bm{x}\in A$, then $11_{A}(\bm{x})=$ 1 and if $\bm{x}\notin A$, $01_{A}(\bm{x})=$ 0.
II-A Problem Statement
This paper considers stochastic discrete-time systems that are modeled by stochastic difference equations of the following form:
$\begin{split}\bm{x}(l+1)=\bm{f}(\bm{x}(l),\bm{\theta}(l)),\forall l\in\mathbb{N},\end{split}$ (1)
where $\bm{x}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ are states and $\bm{\theta}(\cdot)\colon\mathbb{N}\rightarrow\Theta$ with $\Theta\subseteq\mathbb{R}^{m}$ are stochastic disturbances. In addition, let $\bm{\theta}(0),\bm{\theta}(1),\ldots$ be i.i.d. random vectors distributed according to $\mathbb{P}$, with support $\Theta$. The expectation $\mathbb{E}[\cdot]$ is defined with respect to this distribution.
Before defining the trajectory of system (1), we define a disturbance signal.
Definition 1
A disturbance signal $\pi$ is an ordered sequence $\{\bm{\theta}(i),i\in\mathbb{N}\}$, where $\bm{\theta}(\cdot)\colon\mathbb{N}\rightarrow\Theta$.
Given system (1), a signal $\pi=\{\bm{\theta}(i),i\in\mathbb{N}\}$ is a stochastic process defined on the canonical sample space $\Omega=\Theta^{\infty}$, endowed with its product topology $\mathcal{B}(\Omega)$, with the probability measure $\mathbb{P}^{\infty}$. The expectation associated with the probability measure $\mathbb{P}^{\infty}$ is denoted by $\mathbb{E}^{\infty}[\cdot]$.
A disturbance signal $\pi$ together with an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$ induces a unique discrete-time trajectory as follows.
Definition 2
Given a disturbance signal $\pi\in\Omega$ and an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$, a trajectory of system (1) is denoted as $\bm{\phi}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ with $\bm{\phi}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}$, i.e., $\bm{\phi}_{\pi}^{\bm{x}_{0}}(l+1)=\bm{f}(\bm{\phi}_{\pi}^{\bm{x}_{0}}(l),\bm{\theta}(l)),\forall l\in\mathbb{N}.$
In this study, we address two verification problems for the system governed by (1) with a finite-time horizon $[0,N]$, where $N\in\mathbb{N}$. The first problem pertains to a finite-time safety verification problem, examining the likelihood of the system exiting the safe set $\mathcal{X}$ throughout its evolution over $[0,N]$, starting from $\bm{x}_{0}\subseteq\mathcal{X}$. The second one involves a finite-time reach-avoid task, focusing on the probability that the system enters the target set $\mathcal{X}_{r}\subset\mathcal{X}$ safely within the time horizon $[0,N]$, given an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
Problem 1 (Finite-time Safety Verification)
Given a finite time interval $[0,N]$ with $N\in\mathbb{N}$, a safe set $\mathcal{X}$, and an initial state $\bm{x}_{0}\in\mathcal{X}$, the finite-time safety verification problem is to compute $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ which are respectively the lower and upper bounds of the exit probability, with which system (1) starting from $\bm{x}_{0}$ will exit the safe set $\mathcal{X}$ within $[0,N]$, i.e.,
$\epsilon_{1}\leq\mathbb{P}^{\infty}\left(\begin{aligned} \exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\notin\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\right)\leq\epsilon_{2}.$
Problem 2 (Finite-time Reach-avoid Verification)
Given a finite time interval $[0,N]$ with $N\in\mathbb{N}$, a safe set $\mathcal{X}$, a target set $\mathcal{X}_{r}\subseteq\mathcal{X}$, and an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, the finite-time reach-avoid verification problem is to compute $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ which are respectively the lower and upper bounds of the probability, with which system (1) starting from $\bm{x}_{0}$ will enter the target set $\mathcal{X}_{r}$ within $[0,N]$ while staying inside the set $\mathcal{X}$ before the first target hitting time, i.e.,
$\epsilon_{1}\leq\mathbb{P}^{\infty}\left(\begin{aligned} &\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\right)\leq\epsilon_{2}.$
Like stochastic barrier functions based methods, which formulate sufficient conditions for mainly computing upper bounds of the probability of reaching unsafe sets eventually, we in this paper will propose barrier-like conditions for addressing Problem 1 and 2.
It is noteworthy to mention that by treating system (1) within an extended domain $\widehat{\mathcal{X}}$, which includes the safe set $\mathcal{X}$ and all one-step reachable states from $\mathcal{X}$, and interpreting $\widehat{\mathcal{X}}\setminus\mathcal{X}$ as a target region as shown in the sequel, the conditions derived for Problem 2 can be adapted to tackle Problem 1 since $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\notin\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X})=\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\widehat{\mathcal{X}}\mid\bm{x}_{0}\in\mathcal{X})$. However, Problem 1 and 2 are different. From a reachability perspective, the discrepancy arises from whether the target set is contained within the safe set $\mathcal{X}$ or not. When the target set is outside the safe set, as in the safety verification problem in Problem 1, reaching the target set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ is equivalent to exiting the safe set $\mathcal{X}$. However, this does not hold for the reach-avoid problem in Problem 2. Consequently, the conditions derived for Problem 2 should be refined to address Problem 1. Thus, to underscore the improvements and maintain clarity, we treat these problems as separate entities in this paper.
III Finite-time Safety Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds of the probability in Problem 1. The sufficient condition for lower bounds is formulated in Subsection III-A and the one for upper bounds is introduced in Subsection III-B.
In accordance with the methodology in [16], we define a switched system that is constructed by freezing the dynamics of system (1) upon exiting the safe set $\mathcal{X}$. This switched system will facilitate the construction of sufficient conditions for lower and upper bounding the exit probability in the sequel.
Definition 3
The switched stochastic discrete-time system, which is built upon system (1), is a quadruple (~$(\widetilde{\mathcal{L}},\widetilde{\mathcal{X}},\bm{x}_{0},\widetilde{\bm{f}})$ with the following components:
  1. -
    ~=$\widetilde{\mathcal{L}}=\{1,2\}$ is a set of three locations;
  2. -
    $\widetilde{\mathcal{X}}\subseteq\mathbb{R}^{n}$ is the state constraint set;
  3. -
    $\bm{x}_{0}\in\widetilde{\mathcal{X}}$ is the initial state;
  4. -
    $\widetilde{\bm{f}}(\cdot,\cdot)\colon\mathbb{R}^{n}\times\Theta\rightarrow\mathbb{R}^{n}$, where
    $\widetilde{\bm{f}}(\bm{x},\bm{\theta})=\sum_{i=1}^{2}1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})\widetilde{\bm{f}}_{i}(\bm{x},\bm{\theta})$
    with $\widetilde{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$ and $\widetilde{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widetilde{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widetilde{\mathcal{X}}_{i}$, i.e., $11_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=$ 1 if $\bm{x}\in\widetilde{\mathcal{X}}_{i}$; otherwise, $01_{\widetilde{\mathcal{X}}_{i}}(\bm{x})=$ 0,
where
  1. 1.
    $\widetilde{\mathcal{X}}$ is a set satisfying $\widetilde{\Omega}\subset\widetilde{\mathcal{X}}$, where $\widetilde{\Omega}$ is the union of $\mathcal{X}$ and all reachable states starting from $\mathcal{X}$ in one step, i.e.,
    $\widetilde{\Omega}=\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{\theta}),\bm{x}_{0}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
  2. 2.
    $\widetilde{\mathcal{X}}_{1}=\mathcal{X}$;
  3. 3.
    $\widetilde{\mathcal{X}}_{2}=\widetilde{\mathcal{X}}\setminus\mathcal{X}$.
The evolution of the state of this system is governed by the iterative map
$\begin{split}&\bm{x}(l+1)=\widetilde{\bm{f}}(\bm{x}(l),\bm{\theta}(l)),\forall l\in\mathbb{N},\\ &\bm{x}(0)=\bm{x}_{0}\in\widetilde{\mathcal{X}}.\end{split}$ (2)
The trajectory of system (2), induced by initial state $\bm{x}_{0}\in\widetilde{\mathcal{X}}$ and disturbance policy $\pi\in\Omega$, is denoted by $\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
It is observed that if system (1) starting from $\bm{x}_{0}\in\mathcal{X}$ enters the set $\widetilde{\mathcal{X}}\setminus\mathcal{X}$ initially at time $i\in\mathbb{N}$, system (2) starting from $\bm{x}_{0}\in\mathcal{X}$ will also enter it for the first time at this time instant, and vice versa. Furthermore, when system (2) transitions into the set $\widetilde{\mathcal{X}}\setminus\mathcal{X}$, it remains there indefinitely. Thus, the set $\widetilde{\mathcal{X}}$ is a robust invariant for system (2) [16], that is, trajectories of system (2) originating from the set $\widetilde{\mathcal{X}}$ will never leave it. Consequently, the probability of system (2) entering $\widetilde{\mathcal{X}}\setminus\mathcal{X}$ at time $t=i$ is equal to the probability of system (1) entering this set up to time $t=i$.
Lemma 1
For $i\in\mathbb{N}$, $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq i}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X})=\mathbb{P}^{\infty}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X})$. Thus, $\mathbb{P}^{\infty}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X})\leq\mathbb{P}^{\infty}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(j)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X})$ if $i\leq j$.
III-A Sufficient Conditions for Upper Bounds
In this subsection, we present our sufficient condition for upper bounding the probability of exiting the safe set over the time horizon $[0,N]$ in Problem 1.
Theorem 1
If there exist a function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$, and $\alpha\in(0,1]$ and $\beta\in(-\infty,1]$ such that
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (3)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\leq$
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in[0,1]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\gamma\in[0,1]\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha\in(0,1]\wedge\gamma\in(-\infty,0)\end{cases},$
where $\gamma=\beta\alpha-(\alpha-1)$.
Proof:
1) We first prove the case of $\alpha\in(0,1)\wedge\gamma\in[0,1]$.
Since $\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$, we can obtain
$\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-\frac{1}{\alpha}v(\bm{x})=(1-\frac{1}{\alpha})v(\bm{x}),\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}.$
Therefore,
$\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-\frac{1}{\alpha}v(\bm{x})\leq\beta,\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X},$
which can be justified based on the facts that $v(\bm{x})\geq 1$ for $\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$, $\alpha\in(0,1)$ and $\gamma\in[0,1]$. Therefore, if $v(\bm{x})$ satisfies (3), it will satisfy
$\begin{cases}\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\widetilde{\mathcal{X}}\\ v(\bm{x})\geq 1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})&\forall\bm{x}\in\widetilde{\mathcal{X}}\end{cases}.$ (4)
According to (4), we have
$\begin{split}&v(\bm{x}_{0})\geq 1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x}_{0}),\\ &\alpha^{-1}v(\bm{x}_{0})+\beta\geq\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]\geq\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))],\\ &\ldots,\\ &\alpha^{-N}v(\bm{x}_{0})+\beta\sum_{i=0}^{N-1}\alpha^{-i}\geq\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\geq\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))].\end{split}$
Therefore, $\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\leq\alpha^{-N}v(\bm{x}_{0})+\alpha\beta\frac{(1-\alpha^{-N})}{\alpha-1}$. According to Lemma 1, we have the conclusion.
2) The conclusion for the case of $\alpha=1\wedge\gamma\in[0,1]$ can be justified via following the proof of the above one.
3) We will show the case of $\alpha\in(0,1]\wedge\gamma\in(\infty,0)$.
Since $\beta<1-\frac{1}{\alpha}$, thus $\frac{1}{\alpha}<1-\beta$. Consequently, $01-\beta>$ 0. Thus, we conclude that if $v(\bm{x})$ satisfies (3), it will satisfy
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq v(\bm{x})(1-\beta)+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$
and thus it satisfies
$\begin{cases}\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq(1-\beta)v(\bm{x})+\beta&\forall\bm{x}\in\widetilde{\mathcal{X}}\\ v(\bm{x})\geq 1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})&\forall\bm{x}\in\widetilde{\mathcal{X}}\end{cases}.$
Therefore, following the proof for the case $\alpha\in(0,1]\wedge\gamma\in[0,1]$, we will have the conclusion.
The proof is completed. ∎
In Theorem 1, the parameter $\beta$ is practically constrained to be less than or equal to one. This restriction is necessary to ensure the practical utility of the upper bounds $v(\bm{x}_{0})+\beta N$ and $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}$. Assigning a value greater than one to $\beta$ would lead to these two bounds exceeding unity when $N\geq 1$, rendering them ineffective.
If $\alpha=1$ and $\beta=0$, Theorem 1 is equivalent to Proposition 3 in [16], which formulates a sufficient condition for determining an upper bound of the probability of exiting the safe set $\mathcal{X}$ eventually(i.e., $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$). The upper bound is $v(\bm{x}_{0})$, which is also an upper bound of the exit probability $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$. Besides, since $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=\alpha^{-N}(v(\bm{x}_{0})-\frac{\alpha\beta}{\alpha-1})+\frac{\alpha\beta}{\alpha-1}$, we conclude that when $\gamma\in[0,1]$ and $\alpha\in(0,1)$, if $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$, $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$ is also an upper bound of the probability of exiting the safe set $\mathcal{X}$ eventually; if $v(\bm{x}_{0})<\frac{\alpha\beta}{\alpha-1}$,
$\lim_{N\rightarrow\infty}v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=-\infty,$
which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$ and thus $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$. Similarly, when $\alpha\in(0,1)$ and $\gamma\in(-\infty,0)$, if $v(\bm{x}_{0})<1$, $\lim_{N\rightarrow\infty}1-(1-v(\bm{x}_{0}))(1-\beta)^{N}=-\infty$ holds, implying $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$ and $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=0$.
Theorem 1 extends the conclusion drawn in [6], as it allows for a value of $\alpha$ within the interval $(0,1)$, in contrast to the requirement of $\alpha\geq 1$ in the cited work. For convenience of reference, we present the related result in [6] here. If there exists a continuous nonnegative function satisfying
$\begin{cases}v(\bm{x}_{0})<1&\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (5)
$\alpha\geq 1$, and $\beta\geq 0$, then $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\leq$
$\begin{cases}1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha>1\wedge\gamma\in(-\infty,0]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha>1\wedge\gamma\in(0,\infty)\\ v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in(0,\infty)\end{cases},$
where $\gamma=\alpha\beta-(\alpha-1)$.
Since $\bm{x}_{0}\in\{\bm{x}\in\widetilde{\mathcal{X}}\mid v(\bm{x})<1\}$ and $\widetilde{\mathcal{X}}\setminus\mathcal{X}\subseteq\{\bm{x}\in\widetilde{\mathcal{X}}\mid v(\bm{x})\geq 1\}$, according to Theorem 3 in Chapter 3 in [6], the above statement holds. Similarly, the parameter $\beta$ in (5) should be less than or equal to one. When $\gamma\leq 0$ and $\alpha>1$, $\beta\leq 1-\frac{1}{\alpha}$ holds and thus $\beta\leq 1$ holds. For the case with $\gamma>0$, when $\beta>1$ and $N\geq 1$, $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}>1$ and $v(\bm{x}_{0})+\beta N>1$ hold, which are ineffective upper bounds.
When comparing constraints (3) and (5), it becomes apparent that if there exists a function $v(\bm{x})$ along with $\alpha>1$ and $\beta\in[0,1]$ satisfying (5), the same function $v(\bm{x})$ with $\frac{1}{\alpha}$ and $\beta\in[0,1]$ will satisfy (3). Consequently, is can be concluded that (3) is more expressive than (5).
III-B Sufficient Conditions for Lower Bounds
In this subsection, we present our sufficient condition for lower bounding the probability of exiting the safe set over the time horizon $[0,N]$ in Problem 1, which is inspired by [13, 15].
Theorem 2
If there exist a function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ with $\sup_{\bm{x}\in\widetilde{\mathcal{X}}}v(\bm{x})\leq M$, $\alpha\in[1,\infty)$, and $\beta\in(1-\alpha,\infty)$ such that
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (6)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq$
$\begin{cases}\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)}&\text{if~{}}\alpha>1\\ 1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)}&\text{if~{}}\alpha=1\end{cases}.$
Proof:
Since $\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$, we conclude that if $v(\bm{x})$ satisfies (6), it will satisfy
$(\alpha-1+\beta)1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})+\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\geq\alpha v(\bm{x})+\beta,\forall\bm{x}\in\widetilde{\mathcal{X}}.$
Consequently, for $\bm{x}_{0}\in\mathcal{X}$, we have
$\begin{split}&\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]-\alpha v(\bm{x}_{0})\geq\beta+(1-\alpha-\beta)1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\bm{x}_{0}),\\ &\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(2))]-\alpha\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]\geq\beta+(1-\alpha-\beta)\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))],\\ &\ldots,\\ &\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N+1))]-\alpha\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\geq\beta+(1-\alpha-\beta)\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))].\end{split}$
Thus, we obtain
$\begin{split}&\mathbb{E}^{\infty}[v(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N+1))]-\alpha^{N+1}v(\bm{x}_{0})\\ &\geq\beta\sum_{i=0}^{N}\alpha^{i}+(1-\alpha-\beta)\sum_{i=0}^{N}\alpha^{N-i}\mathbb{E}^{\infty}[1_{\widetilde{\mathcal{X}}\setminus\mathcal{X}}(\widetilde{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i))]\\ &\geq\beta\frac{1-\alpha^{N+1}}{1-\alpha}+(1-\alpha-\beta)\frac{1-\alpha^{N+1}}{1-\alpha}\times\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\end{split}.$
The last inequality is obtained according to Lemma 1, which states $\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}$ for $i\leq N$, and the fact that $01-\alpha-\beta<$ 0.
Consequently,
$\begin{split}&\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)}\end{split}.$
The proof is completed. ∎
It is worth noting here that if there exists a bounded function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying (6) with $\alpha=1$, then system (1), starting from $\bm{x}_{0}\in\mathcal{X}$, will exit the safe set $\mathcal{X}$ eventually with the probability of one, since $\lim_{N\rightarrow\mathbb{N}}1+\frac{v(\bm{x}_{0})-M}{\beta(N+1)}=1$. Therefore, if system (1) does not feature this property, $\alpha=1$ cannot be used to perform computations.
Remark 1
When $\alpha=1$, the constraint $v(\bm{x})\leq 1,\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}$ in (6) is redundant and can be removed.
Remark 2
One may wander whether a sufficient condition for lower bounding the probability of exiting the safe set $\mathcal{X}$ over the time horizon $[0,N]$ in Problem 1 can be constructed by directly reversing the sign in constraint (3) in Theorem 1. We will give a brief explanation here that the condition constructed in this manner will consistently yield zero as lower bounds.
If there exists a function $v(\bm{x})\colon\widetilde{\mathcal{X}}\rightarrow\mathbb{R}$ such that
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\geq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (7)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=\mathbb{P}^{\infty}\big{(}\widetilde{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\widetilde{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq$
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\gamma\in(-\infty,0)\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\gamma\in(-\infty,0)\\ 1-(1-v(\bm{x}_{0}))\alpha^{-N}&\text{if~{}}\alpha\in(0,1]\wedge\gamma\in[0,\infty)\end{cases},$
where $\gamma=\beta\alpha-(\alpha-1)$. This conclusion can be justified by following the proof of Theorem 1. However, via $\gamma=\beta\alpha-(\alpha-1)<0$ and $\alpha\in(0,1]$, we have $\beta<0$. Also, since $v(\bm{x})\leq 0$ for $\bm{x}\in\mathcal{X}$, $v(\bm{x}_{0})\leq 0$ holds. Thus, $v(\bm{x}_{0})+\beta N\leq 0$, $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}\leq 0$, and $1-(1-v(\bm{x}_{0}))\alpha^{-N}\leq 0$ hold.
IV Finite-time Reach-avoid Verification
In this section, we present our sufficient conditions for characterizing upper and lower bounds of the probability in Problem 2. The sufficient condition for upper bounds is formulated in Subsection IV-A and the one for lower bounds is introduced in Subsection IV-B.
Similar to [13], we define a switched system, which facilitates the transformation of the reach-avoid problem in Problem 2 to a mere reachability problem. The switched system is constructed by freezing the dynamics of system (1) upon either exiting the safe set $\mathcal{X}$ or reaching the target set $\mathcal{X}_{r}$.
Definition 4
The switched stochastic discrete-time system, which is built upon system (1), is a quadruple (^$(\widehat{\mathcal{L}},\widehat{\mathcal{X}},\bm{x}_{0},\widehat{\bm{f}})$ with the following components:
  1. -
    ^=$\widehat{\mathcal{L}}=\{1,2,3\}$ is a set of three locations;
  2. -
    $\widehat{\mathcal{X}}\subseteq\mathbb{R}^{n}$ is the state constraint set;
  3. -
    $\bm{x}_{0}\in\widehat{\mathcal{X}}$ is the initial state;
  4. -
    $\widehat{\bm{f}}(\cdot,\cdot)\colon\mathbb{R}^{n}\times\Theta\rightarrow\mathbb{R}^{n}$, where
    $\widehat{\bm{f}}(\bm{x},\bm{\theta})=\sum_{i=1}^{3}1_{\widehat{\mathcal{X}}_{i}}(\bm{x})\widehat{\bm{f}}_{i}(\bm{x},\bm{\theta})$
    with $\widehat{\bm{f}}_{1}(\bm{x},\bm{\theta})=\bm{f}(\bm{x},\bm{\theta})$, $\widehat{\bm{f}}_{2}(\bm{x},\bm{\theta})=\bm{x}$ and $\widehat{\bm{f}}_{3}(\bm{x},\bm{\theta})=\bm{x}$, and $1_{\widehat{\mathcal{X}}_{i}}(\bm{x})$ is the indicator function of the set $\widehat{\mathcal{X}}_{i}$, i.e., $11_{\widehat{\mathcal{X}}_{i}}(\bm{x})=$ 1 if $\bm{x}\in\widehat{\mathcal{X}}_{i}$; otherwise, $01_{\widehat{\mathcal{X}}_{i}}(\bm{x})=$ 0,
where
  1. 1.
    $\widehat{\mathcal{X}}$ is a set satisfying $\widehat{\Omega}\subset\widehat{\mathcal{X}}$, where
    $\widehat{\Omega}=\{\bm{x}\in\mathbb{R}^{n}\mid\bm{x}=\bm{f}(\bm{x}_{0},\bm{\theta}),\bm{x}_{0}\in\mathcal{X},\bm{\theta}\in\Theta\}\cup\mathcal{X};$
  2. 2.
    $\widehat{\mathcal{X}}_{1}=\mathcal{X}\setminus\mathcal{X}_{r}$;
  3. 3.
    $\widehat{\mathcal{X}}_{2}=\mathcal{X}_{r}$;
  4. 4.
    $\widehat{\mathcal{X}}_{3}=\widehat{\mathcal{X}}\setminus\mathcal{X}$.
The evolution of the state of this switched system is governed by the iterative map
$\begin{split}&\bm{x}(l+1)=\widehat{\bm{f}}(\bm{x}(l),\bm{\theta}(l)),\forall l\in\mathbb{N},\\ &\bm{x}(0)=\bm{x}_{0}\in\widehat{\mathcal{X}}.\end{split}$ (8)
The trajectory of system (8), induced by initial state $\bm{x}_{0}\in\widehat{\mathcal{X}}$ and disturbance policy $\pi\in\Omega$, is denoted by $\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$.
When system (1), initialized at $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, enters the region $\mathcal{X}_{r}$ for the first time at time $i\in\mathbb{N}$ without leaving the safe set $\mathcal{X}$ beforehand, system (8) starting from $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$ also experiences its initial entry at this time, and vice versa. Further, when entering $\mathcal{X}_{r}$ and $\widehat{\mathcal{X}}\setminus\mathcal{X}$, system (8) will remain confined indefinitely. Thus, the set $\widehat{\mathcal{X}}$ is a robust invariant for system (8) [13], that is, trajectories of system (8) originating from the set $\widehat{\mathcal{X}}$ will never leave it. Consequently, the probability of system (8) entering $\mathcal{X}_{r}$ at time $t=i$ aligns with the cumulative probability of system (1) entering this target set up to time $i$ without exiting $\mathcal{X}$ until the target hitting event occurs.
Lemma 2
For $i\in\mathbb{N}$, $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq i}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})=\mathbb{P}^{\infty}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})$. Thus, $\mathbb{P}^{\infty}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})\leq\mathbb{P}^{\infty}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(j)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})$ if $i\leq j$.
IV-A Sufficient Conditions for Upper Bounds
In this subsection, we introduce our sufficient condition for upper bounding the probability in Problem 2.
Theorem 3
If there exist a function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$, $\alpha\in(0,1]$, and $\beta\in[0,1]$ such that
$\begin{cases}\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}_{r}\end{cases},$ (9)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\leq$
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\wedge\beta\in[0,1]\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha\in(0,1)\wedge\beta\in[0,1]\end{cases}.$
Proof:
1) We first justify the case of $\alpha\in(0,1)\wedge\beta\in[0,1]$.
Since $\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}$, we can obtain
$\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-\frac{1}{\alpha}v(\bm{x})=(1-\frac{1}{\alpha})v(\bm{x}),\forall\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}.$
Therefore,
$\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]-\frac{1}{\alpha}v(\bm{x})\leq 0\leq\beta,\forall\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r},$
which can be justified based on the facts that $v(\bm{x})\geq 0$ for $\bm{x}\in\widehat{\mathcal{X}}$, $\alpha\in(0,1)$ and $\beta\in[0,1]$. Therefore, if $v(\bm{x})$ satisfies (9), it will satisfy
$\begin{cases}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\widehat{\mathcal{X}}\\ v(\bm{x})\geq 1_{\mathcal{X}_{r}}(\bm{x})&\forall\bm{x}\in\widehat{\mathcal{X}}\end{cases}.$ (10)
According to (10), we have
$\begin{split}&v(\bm{x}_{0})\geq 1_{\mathcal{X}_{r}}(\bm{x}_{0}),\\ &\alpha^{-1}v(\bm{x}_{0})+\beta\geq\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]\geq\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))],\\ &\ldots,\\ &\alpha^{-N}v(\bm{x}_{0})+\beta\sum_{i=0}^{N-1}\alpha^{-i}\geq\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\geq\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))].\end{split}$
Thus, $\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\leq\alpha^{-N}v(\bm{x}_{0})+\alpha\beta\frac{(1-\alpha^{-N})}{\alpha-1}.$ According to Lemma 2, we have the conclusion.
2) The conclusion on the case of $\alpha=1\wedge\beta\in[0,1]$ can be justified by following the proof for the above one.
The proof is completed. ∎
Similar to the analysis for the parameter $\beta$ in Theorem 1, the parameter $\beta$ in Theorem 3 is also required to be less than or equal to 1, which is subject to a practical restriction. Assigning a value greater than one to $\beta$ would result in bounds exceeding unity when $N\geq 1$, rendering them ineffective.
On the other hand, the analysis for the upper bounds in Theorem 3 is similar to the one in Theorem 1. When $\alpha=1$ and $\beta=0$, the upper bound is $v(\bm{x}_{0})$, which is an upper bound of the probability $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}$. Besides, since $v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=\alpha^{-N}(v(\bm{x}_{0})-\frac{\alpha\beta}{\alpha-1})+\frac{\alpha\beta}{\alpha-1}$, we conclude that when $\alpha\in(0,1)$ and $\beta\in[0,1]$, if $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$, $v(\bm{x}_{0})=\frac{\alpha\beta}{\alpha-1}$ is also an upper bound of the probability of reaching the target set $\mathcal{X}_{r}$ eventually while staying within the safe set $\mathcal{X}$ before the first target hitting time; if $v(\bm{x}_{0})<\frac{\alpha\beta}{\alpha-1}$, $\lim_{N\rightarrow\infty}v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}=-\infty$, which implies $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=0$ and thus $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=0$.
Remark 3
In Theorem 3, $\alpha$ is limited to be in $(0,1]$. We can obtain upper bounds for the case with $\alpha>1$ according to Theorem 3 in Chapter 3 in [6], as shown in Theorem 4.
Theorem 4
If there exist a continuous non-negative function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$, $\alpha\in[1,\infty)$, and $\beta\in[0,1]$ such that
$\begin{cases}v(\bm{x})<1&\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\leq\frac{v(\bm{x})}{\alpha}+\beta&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{r}\cup(\widehat{\mathcal{X}}\setminus\mathcal{X})\\ v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\end{cases},$ (11)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\leq$
$\begin{cases}v(\bm{x}_{0})+\beta N&\text{if~{}}\alpha=1\\ v(\bm{x}_{0})\alpha^{-N}+\frac{(1-\alpha^{-N})\alpha\beta}{\alpha-1}&\text{if~{}}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}>1\\ 1-(1-v(\bm{x}_{0}))(1-\beta)^{N}&\text{if~{}}\alpha>1\wedge\frac{\beta\alpha}{\alpha-1}\leq 1\end{cases}.$ (12)
Proof:
Since $v(\bm{x})\geq 1$ for $\bm{x}\in\mathcal{X}_{r}\cup(\widehat{\mathcal{X}}\setminus\mathcal{X})$, we have
$\begin{split}&\mathbb{P}^{\infty}\left(\begin{aligned} &\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}_{r}\wedge\\ &\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(l)\in\mathcal{X}\end{aligned}\middle|\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\right)\leq\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.v(\bm{\phi}^{\bm{x}_{0}}_{\pi}(k))\geq 1\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}).\end{split}$
Also, since $\bm{x}_{0}\in\{\bm{x}\in\widehat{\mathcal{X}}\mid v(\bm{x})<1\}$, Theorem 3 in Chapter 3 in [6] implies that the bounds in (12) are upper bounds for $\mathbb{P}^{\infty}(\exists k\in\mathbb{N}_{\leq N}.v(\bm{\phi}^{\bm{x}_{0}}_{\pi}(k))\geq 1\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r})$ and thus the conclusion holds. ∎
Following the comparison between constraints (3) and (5), we can conclude that constraint (9) is more expressive that constraint (11).
Remark 4
Inspired by Theorem 3 in Chapter 3 in [6], $c$-martingale for stochastic discrete-time systems was proposed for temporal verification in [5]. Under the assumption that $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ (i.e., $\mathcal{X}$ is a robust invariant set for system (1)), the $c$-martingale in [5] is to certify upper bounds of the probability with which system (1) starting from $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$ will reach the set $\mathcal{X}_{r}$ within a specified bounded time horizon. Under this strong assumption, $c$-martingale satisfies (9) in Theorem 3 with $\alpha=1$. This assumption is also imposed in [17]. It is worth noting here that $\widehat{\mathcal{X}}$ in our conditions is not needed if the assumption that $\bm{f}(\cdot,\cdot)\colon\mathcal{X}\times\Theta\rightarrow\mathcal{X}$ holds.
IV-B Sufficient Conditions for Lower Bounds
In this subsection, we introduce our sufficient condition for lower bounding the probability in Definition 2.
Theorem 5
If there exist a function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$ with $\sup_{\bm{x}\in\widehat{\mathcal{X}}}v(\bm{x})\leq M$, $\alpha\in(1,\infty)$, and $\beta\in(1-\alpha,\infty)$ such that
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ (\alpha-1)v(\bm{x})\leq-\beta&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (13)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}_{\leq N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\bigwedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}=\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\geq$
$\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)}.$
Proof:
Since $\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}$, we can obtain that if $v(\bm{x})$ satisfies (13), it will satisfy
$(\alpha-1+\beta)1_{\mathcal{X}_{r}}(\bm{x})+\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\geq\alpha v(\bm{x})+\beta,\forall\bm{x}\in\widehat{\mathcal{X}}.$ (14)
From (14), for $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we have
$\begin{split}&\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]-\alpha v(\bm{x}_{0})\geq\beta+(1-\alpha-\beta)1_{\mathcal{X}_{r}}(\bm{x}_{0}),\\ &\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(2))]-\alpha\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))]\geq\beta+(1-\alpha-\beta)\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(1))],\\ &\ldots,\\ &\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N+1))]-\alpha\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))]\geq\beta+(1-\alpha-\beta)\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N))].\end{split}$
Thus, we can obtain
$\begin{split}&\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(N+1))]-\alpha^{N+1}v(\bm{x}_{0})\\ &\geq\beta\sum_{i=0}^{N}\alpha^{i}+(1-\alpha-\beta)\sum_{i=0}^{N}\alpha^{N-i}\mathbb{E}^{\infty}[1_{\mathcal{X}_{r}}(\widehat{\bm{\phi}}_{\pi}^{\bm{x}_{0}}(i))]\\ &\geq\beta\frac{1-\alpha^{N+1}}{1-\alpha}+(1-\alpha-\beta)\frac{1-\alpha^{N+1}}{1-\alpha}\times\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\end{split}.$
The last inequality is obtained via Lemma 2, which states $\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\geq\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}$ for $i\leq N$, and the fact that $01-\alpha-\beta<$ 0.
Consequently,
$\begin{split}&\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\geq\frac{(\alpha^{N+1}v(\bm{x}_{0})-M)(\alpha-1)+\beta(\alpha^{N+1}-1)}{(\alpha+\beta-1)(\alpha^{N+1}-1)}\end{split}.$
According to Lemma 2, we have the conclusion. The proof is completed. ∎
Remark 5
Comparing Theorem 2 and 5, we observe that $\alpha$ cannot be equal to one in Theorem 5. Since if $\alpha=1$, we have $\beta>0$ from $\beta\in(1-\alpha,\infty)$. However, we have $\beta\leq 0$ from $(\alpha-1)v(\bm{x})\leq-\beta,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$. This is a contradiction.
V Examples
In this section, we demonstrate the performance of the proposed conditions for safety and reach-avoid verification on two numerical examples, using the semi-definite programming tool Mosek [3].
TABLE I: Computed upper bounds of the probability in Problem 1 in Example 1
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(5) with $\alpha=1.1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.9795 0.9435 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427
(3) with $\alpha=\frac{1}{1.1}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8166 0.1564 0.0650 0.0447 0.0404 0.0398 0.0398 0.0398 0.0398 0.0398
(3) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351
TABLE II: Computed upper bounds of the probability in Problem 2 in Example 1
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(11) with $\alpha=1.1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 0.9868 0.9441 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427
(9) with $\alpha=\frac{1}{1.1}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 0.2530 0.1260 0.0970 0.0906 0.0898 0.0897 0.0897 0.0897 0.0897
(9) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736
Example 1
Consider the following one-dimensional discrete-time system:
$x(l+1)=x(l)+d(l),$
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-0.1,0.1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $x_{0}=0.2$, $\mathcal{X}_{r}=\{\,x\mid(x-0.9)^{2}-10^{-4}\leq 0\,\}$, and $N=30$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods are around 0.0085 and 0.0128, respectively.
Given the small exact probabilities in Problems 1 and 2, we only estimate their upper bounds in this example. The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{x\mid x^{2}-2\leq 0\}$ is employed in solving (3), (5), (9), and (11). To address these constraints, we encode them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed upper bounds are summarized in Tables I and II. Table I illustrates that the constraint (3) with $\alpha=\frac{1}{1.1}$ can yield tighter upper bounds for the probability in Problem 1 compared to (3) with $\alpha=1$ and (5) with $\alpha=1.1$. Similarly, Table II shows that the constraint (9) with $\alpha=\frac{1}{1.1}$ outperforms (9) with $\alpha=1$ and (11) with $\alpha=1.1$ in providing tighter bounds for the probability in Problem 2. Generally, employing higher-degree polynomials for computations facilitates the gain of tighter upper bounds of the probabilities in Problem 1 and 2. However, it is observed that the upper bounds converge rapidly as the polynomial degree increases for this example.
TABLE III: Computed lower and upper bounds of the probability in Problem 1 in Example 2
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(5) with $\alpha=1.01$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8798 0.7694 0.7423 0.7302 0.6923 0.6663 0.6130 0.6127 0.5837 0.5830
(5) with $\alpha=1.001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8169 0.6612 0.6041 0.5845 0.5316 0.4942 0.4131 0.4098 0.3615 0.3605
(3) with $\alpha=\frac{1}{1.01}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 1.0000 0.9625 0.9235 0.8352 0.7731 0.6342 0.6236 0.5424 0.5397
(3) with $\alpha=\frac{1}{1.001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8515 0.6877 0.6179 0.5929 0.5381 0.4983 0.4097 0.4027 0.3505 0.3488
(3) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8100 0.6542 0.5880 0.5643 0.5123 0.4745 0.3902 0.3835 0.3345 0.3322
(6) with $\alpha=1.1$ and $\beta=0$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{1}$ $2.5096\times 10^{-14}$ 0.0279 0.0490 0.0502 0.0711 0.0735 0.0990 0.1064 0.1245 0.1289
TABLE IV: Computed lower and upper bounds of the probability in Problem 2 in Example 2
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(11) with $\alpha=1.001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(11) with $\alpha=1.0001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(11) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(9) with $\alpha=\frac{1}{1.001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9929 0.9827 0.9717 0.9524 0.9380 0.9208 0.9043 0.8932 0.8858
(9) with $\alpha=\frac{1}{1.0001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9506 0.9403 0.9304 0.9117 0.8979 0.8816 0.8658 0.8550 0.8480
(9) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9460 0.9357 0.9259 0.9073 0.8936 0.8774 0.8617 0.8508 0.8439
(13) with $\alpha=1.06$ and $\beta=0$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{1}$ 0.1591 0.2824 0.3453 0.3669 0.4606 0.5218 0.5732 0.5778 0.6119 0.6128
Example 2
Consider the following one-dimensional discrete-time system from [16]:
$x(l+1)=(-0.5+d(l))x(l),$
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-1,1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $x_{0}=-0.9$, $\mathcal{X}_{r}=\{\,x\mid x^{2}-0.36\leq 0\,\}$, and $N=50$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods are around 0.2321 and 0.7708, respectively.
The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{\,x\mid x^{2}-2.25\leq 0\,\}$ is used in solving (3), (5), (6), (9), (11), and (13). These constraints are addressed via encoding them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed lower and upper bounds are summarized in Tables III and IV. It is evident from the results that employing higher-degree polynomials for computations leads to tighter lower and upper bounds of the probabilities in Problems 1 and 2, and the constraints (3) and (5) can complement each other in providing upper bounds of the probability in Problem 1. However, the performance of the constraint (9), with $\alpha=\frac{1}{1.001}$, $\alpha=\frac{1}{1.0001}$, and $\alpha=1$, surpasses that of (11), with $\alpha=1.001$, $\alpha=1.0001$, and $\alpha=1$, in yielding tighter upper bounds of the probability in Problem 2.
In Example 1 and 2, we initially determine upper bounds by assigning pre-defined values to the parameter $\alpha$ in constraints (3), (5), (9) and (11), and lower bounds by assigning pre-defined values to the parameters $\alpha$ and $\beta$ in (6) and (13), and employing convex optimization to solve them. This might yield conservative bounds. However, the automatic optimization of the function $v(\bm{x})$, $\alpha$, and $\beta$ to enhance these bounds constitutes a non-convex problem, which poses a significant challenge. Future work will address this issue, as it is beyond the scope of the current study.
VI Conclusion
In this paper, we introduced novel sufficient conditions for the finite-time safety and reach-avoid verification of stochastic discrete-time dynamical systems. These conditions effectively determine the lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). They complement existing criteria or bridge existing gaps in the literature. Finally, we demonstrated their performance in finite-time safety and reach-avoid verification on two numerical examples, utilizing semi-definite programming tools.